2.5. If G is a finite group, show that there exists a positive integer m such that am = e for all a ∈ G.


1
Expert's answer
2022-06-20T17:54:11-0400

Let "G" be finite group and "1\ne a\in G".

Consider the set "a,a^2,a^3,\dots,a^k,\dots". It is clear that "a^i\ne a^{i+1}" for some integers from the beginning. Since "G" is a finite group there exists "i" and "j" such that "a^i=a^j" implies "a^{i-j}=1". Therefore every element has the finite order. That is the smallest positive integer "k" satisfying "a^k=1". (One may assume without loss of generality that "i>j"). One can do this for each "a\in G". The least common multiple "m" of the order of all elements of "G" satisfies "a^m=1" for all "a\in G".